OpenStudy (anonymous):

10 fourth root of 6 end root plus 4 cubed root of 6 end root minus 9 cubed root of 6 end root minus 3 fourth root of 6

OpenStudy (mathstudent55):

You can combine together the fourth roots. You can combine together the third roots. But you can't combine together the third and fourth roots.

OpenStudy (anonymous):

so the answer would be?

OpenStudy (mathstudent55):

\[10\sqrt[4]{6} + 4\sqrt[3]{6} - 9\sqrt[3]{6} - 3\sqrt[4]{6}\]

OpenStudy (mathstudent55):

is that the problem?

OpenStudy (mathstudent55):

\[10\sqrt[4]{6} - 3\sqrt[4]{6} + 4\sqrt[3]{6} - 9\sqrt[3]{6}\]

OpenStudy (mathstudent55):

\[7\sqrt[4]{6} - 5\sqrt[3]{6}\] Final answer

OpenStudy (anonymous):

thanks& keep the help coming

OpenStudy (mathstudent55):

yw

OpenStudy (anonymous):

got a new problem u should check it out